OpenStudy (anonymous):

Write a quadratic equation having the given numbers as solutions. -3 and-1

jimthompson5910 (jim_thompson5910):

x=-3 or x=-1 x+3 = 0 or x+1 = 0 (x+3)(x+1) = 0 x^2+1x+3x+3 = 0 x^2+4x+3 = 0 So the answer is \[\large x^2+4x+3=0\]